# Eka-Eval : A Comprehensive Evaluation Framework for Large Language Models in Indian Languages

### v1

@@ 4.1.4 Results Processing System
The Results Processing System handles comprehensive output management through three secondary components:
Metrics Calculator:
Computes evaluation metrics and statistical significance using HuggingFace’s evaluate library (e.g., accuracy, BLEU, F1, exact match, Pass@1). It also implements robust error handling for edge cases and missing data, for example, when a model returns “The answer is probably B”, regex-based extraction retrieves the label; if that fails, a default score is assigned. Code completions are sandboxed with timeout control to ensure safe execution and error logging.
Visualisations analytics:
Provides rich analytics including confidence intervals, comparative analysis across multiple models and benchmark configurations. Multiple visualisations like bar charts, heatmaps and radar plots can be generated including support for cross-model performance comparisons.
Export Manager:
Handles result export in multiple formats including JSON, CSV reports to support analysis and visualization workflows. The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations to ensure reproducibility across different experimental setups.
@@ 4.2 Comprehensive Benchmark Coverage
Eka-Eval covers eight major evaluation categories with comprehensive benchmark support across both English and Indic languages:
General Reasoning and Knowledge: For foundational capabilities, we include MMLU Hendrycks et al. (2020) and MMLU-Pro Wang et al. (2024) for multitask language understanding, IFEval Zhou et al. (2023) for instruction following, BBH (3-shot) Suzgun et al. (2022) for challenging reasoning tasks, and AGI-Eval (3-5 shot) Zhong et al. (2023) for general intelligence assessment.
Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K Cobbe et al. (2021) for grade school math, MATH Hendrycks et al. (2021) for competition-level problems, and ARC-Challenge Clark et al. (2018) for scientific reasoning.
Code Generation and Programming: Programming abilities are assessed using HumanEval Chen et al. (2021), MBPP Austin et al. (2021), HumanEval+ Liu et al. (2023), MBPP EvalPlus Liu et al. (2023) with Pass@1 accuracy metrics.
Commonsense Reasoning: We incorporate PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), ARC Easy and Challenge Clark et al. (2018), WinoGrande Sakaguchi et al. (2021), CommonSenseQA Talmor et al. (2018), and OpenBookQA Mihaylov et al. (2018) for comprehensive commonsense evaluation.
World Knowledge: Factual knowledge is tested through TriviaQA (5-shot) Joshi et al. (2017) and NaturalQuestions (5-shot) Kwiatkowski et al. (2019) with accuracy metrics.
@@
Long-Context Understanding: For extended context reasoning, we include ZeroSCROLLS Shaham et al. (2023) with ROUGE and F1 metrics, Needle-in-a-Haystack for retrieval accuracy, and InfiniteBench Zhang et al. (2024) for task-specific long-context evaluation.
@@
4.3 Multilingual and Indic Language Support
A distinguishing feature of Eka-Eval is its comprehensive support for Indic languages, addressing a critical gap in existing evaluation frameworks. The system includes dedicated benchmarks for major Indian languages including Hindi, Bengali, Tamil, Telugu, Marathi, Gujarati, Kannada, Malayalam, Punjabi, and Odia. For multilingual evaluation, we support popular Indic language benchmarks including MMLU-IN, TriviaQA-IN, MILU Verma et al. (2024), GSM8K-IN, BoolQ-IN, ARC-C-IN, Flores-IN with BLEU (Papineni et al., 2002) and ChrF metrics, XQuAD-IN and XorQA-IN with F1 and exact match scores.
The multilingual architecture includes language-specific prompt templates, culturally-appropriate evaluation protocols, and specialized tokenization handling for Indic scripts. This ensures fair and accurate evaluation across linguistic boundaries while maintaining consistency with global benchmarks.
@@ 4.4 Extensibility and Customization
The framework is designed with extensibility as a first-class concern. New benchmarks can be integrated through a simple plugin architecture that requires minimal boilerplate code. The system supports custom evaluation metrics, prompt templates, and post-processing pipelines, enabling researchers to adapt the framework to their specific requirements.
Configuration management is handled through a hierarchical system supporting JSON formats, allowing users to define evaluation suites ranging from quick smoke tests to comprehensive benchmark runs. The configuration system supports parameter sweeping, enabling systematic exploration of prompt variations, few-shot examples, and model hyperparameters.
@@ 5 Conclusion and Future Work
In this work, we introduced Eka-Eval, a unified and extensible framework for the evaluation of large language models across both global and Indic tasks. By supporting a diverse set of benchmarks and accommodating multiple model backends—including local checkpoints and API-based deployments—Eka-Eval enables reproducible, scalable, and backend-agnostic evaluation pipelines.
Looking ahead, we plan to expand the framework to support over 100 tasks, with a particular emphasis on underrepresented Indic languages and culturally grounded evaluations. Future extensions will include categorization of benchmarks by language diversity, task difficulty, and professional domains such as law, healthcare, and governance. We aim to introduce dynamic task calibration mechanisms that vary context length, ambiguity, and reasoning complexity, as well as India-specific knowledge grounding through tasks on government schemes and public health challenges. Additional priorities include designing benchmarks for bias detection, hallucination analysis, and privacy risk assessment; incorporating human-centered meta-evaluation; supporting domain-specific evaluation; and enabling adversarial human review. Finally, we plan to support phase-wise evaluation—spanning pretraining, fine-tuning, and deployment stages—to offer a comprehensive view of model behavior across the entire lifecycle.
These enhancements will be included in Eka-Eval v2.0, strengthening its role as the evaluation standard for Indic and other LLMs.
@@ A.1 Indic benchmark configuration
This figure showcases a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Indic languages. It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
@@
A.2 Demonstration
The interactive CLI of the Eka-Eval framework is shown below, which guides users through model selection and evaluation setup. It simplifies benchmarking workflows, making it accessible to both researchers and developers.
@@ A.2 Demonstration
Users are prompted to select high-level task groups (e.g., Reading Comprehension) during CLI setup. This enables fine-grained benchmarking organization and streamlined selection.
@@
Eka-Eval supports local HuggingFace models and API-based models like OpenAI, Claude, and Gemini. Users interactively select model source and configuration through CLI.
@@
After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain.
@@
The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV.
@@
The framework allows users to generate multiple types of visualizations : bar charts, heatmaps, radar plots that are based on completed evaluations.
@@
This chart shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
@@ A.2.1 Prompt Template System
A critical component of Eka-Eval is its sophisticated prompt management system that handles diverse evaluation paradigms and languages. The framework implements a flexible template system demonstrated through PIQA benchmark prompt 9:
@@
The template system supports zero-shot, few-shot, and chain-of-thought prompting strategies, ensuring consistency across evaluation modes and languages. Users can customize prompt strategies and easily configure them in the benchmark configuration file, as shown in Figure 2.
@@ Appendix B Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it currently lacks a graphical user interface, relying instead on CLI-based workflows. The framework does not yet support vLLMs and provides limited support for detailed error analysis or explainability. Additionally, reproducibility may be affected by changes in external datasets or model versions unless explicitly versioned or cached.

### v2

@@ 4.2.4 Results Processing System
Handles comprehensive output management through three secondary components:
Metrics Calculator: Computes evaluation metrics using HuggingFace’s evaluate library (e.g., accuracy, BLEU, F1-score, exact match, Pass@1). It also implements robust error handling for edge cases and missing data; for example, when a model returns “The answer is probably B”, regex-based extraction retrieves the label; if that fails, a default score is assigned. Code completions are sandboxed with timeout control to ensure safe execution and error logging.
Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps and radar plots (including support for cross-model performance comparisons).
Export Manager: Handles result export in multiple formats including JSON and CSV. The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations.
@@ 4.3 Comprehensive Benchmark Coverage
Eka-Eval covers nine major evaluation categories with comprehensive benchmark support across 35+ benchmarks (See Appendix A). These categories include:
  1. 1.
    Code Generation and Programming: Programming abilities are assessed using HumanEval (Chen et al., 2021), MBPP  (Austin et al., 2021), HumanEval+ (Liu et al., 2023), EvalPlus (Liu et al., 2023), with Pass@1 accuracy metrics.
  2. 2.
    Mathematics and Logical Reasoning: Mathematical capabilities are evaluated through GSM8K (Cobbe et al., 2021) for grade school math, MATH (Hendrycks et al., 2021) for competition-level problems, and ARC-C (Clark et al., 2018) for scientific reasoning.
  3. 3.
@@
  4. 4.
    Commonsense Reasoning: We incorporate PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), ARC-C (Clark et al., 2018), WinoGrande (Sakaguchi et al., 2021), CommonSenseQA (Talmor et al., 2018), and OpenBookQA (Mihaylov et al., 2018) for comprehensive commonsense evaluation.
  5. 5.
    World Knowledge: Factual knowledge is tested through TriviaQA (5-shot) (Joshi et al., 2017) and NaturalQuestions (5-shot) (Kwiatkowski et al., 2019) with accuracy metrics.
  6. 6.
    Long-Context Understanding: For extended context reasoning, we include ZeroSCROLLS (Shaham et al., 2023) with ROUGE (Lin, 2004) and F1 scores, Needle-in-a-Haystack (Wang et al., 2024a) for retrieval accuracy, and InfiniteBench (Zhang et al., 2024) for task-specific long-context evaluation.
  7. 7.
    General Reasoning and Knowledge: For foundational capabilities, we include MMLU (Hendrycks et al., 2020a) and MMLU-Pro (Wang et al., 2024b) for multitask language understanding, IFEval (Zhou et al., 2023) for instruction following, BBH (3-shot) (Suzgun et al., 2022) for challenging reasoning tasks, and AGI-Eval (3-5 shot) (Zhong et al., 2023) for general intelligence assessment.
  8. 8.
@@
  9. 9.
    Multilingual and Indic Language Support: A distinguishing feature of Eka-Eval is its dedicated support for Indian languages, addressing a critical gap in existing evaluation frameworks. The system includes benchmarks covering the majority of Indian languages. For multilingual evaluation, it supports popular Indic language benchmarks including MMLU-IN (Hendrycks et al., 2020b), TriviaQA-IN (Joshi et al., 2017), MILU (Verma et al., 2024), GSM8K-IN (Cobbe et al., 2021), BoolQ-IN (Clark et al., 2019), ARC-C-IN (Clark et al., 2018), Flores-IN (Goyal et al., 2022) with BLEU (Papineni et al., 2002) and ChrF metrics (Popović, 2015), XQuAD-IN (Artetxe et al., 2019) and XorQA-IN (Asai et al., 2021) with F1 and exact match scores.
    The multilingual architecture includes language-specific prompts, culturally appropriate protocols, and specialized tokenization for Indic scripts, ensuring fair evaluation across languages while staying consistent with global benchmarks.
@@ 4.4 Extensibility and Customization
The framework is designed with extensibility as a first-class concern. New benchmarks can be integrated through a simple plugin architecture that requires minimal boilerplate code. The system supports custom evaluation metrics, prompt templates, and post-processing pipelines, enabling users to adapt the framework to their specific requirements.
Configuration is managed through a hierarchical JSON-based system that enables users to define evaluation suites ranging from quick smoke tests to comprehensive benchmark runs. It also supports parameter sweeps, facilitating systematic exploration of prompt variations, few-shot examples, and model hyperparameters.
5 Experiments
To assess the effectiveness and usability of Eka-Eval, we conducted a comprehensive evaluation combining benchmark coverage analysis and user-centered feedback across six prominent frameworks: lm-eval-harness, OpenCompass, HELM, FreeEval, indic-eval, and Eka-Eval.
5.1 Evaluation Procedure
Three human participants independently evaluated each framework by running the Sarvam-1B Sarvam (2024) model on four diverse benchmarks: WinoGrande, PIQA, ARC-C and FLORES. Participants installed each framework in a standardized environment, integrated the target model, and ran the recommended benchmark workflows. They then rated the framework on five criteria using Likert scale ranging from 1 (Very Poor) to 5 (Excellent):
EKA-EVAL Analysis: All ratings were recorded to enable direct comparison across frameworks. Detailed instructions and a standardized rating form were provided to ensure consistency and reproducibility. Participants were encouraged to consult official documentation but not to seek external technical support beyond publicly available resources.
Participants assessed each framework across five evaluation criteria:
  1. 1.
    Setup and Configuration Time: Time and effort required to install dependencies, configure models, and execute an initial benchmark run.
  2. 2.
    Ease of Navigation: Intuitiveness of navigation, benchmark selection, and configuration, including CLI clarity, documentation quality, and ease of discovering options.
  3. 3.
    Result Reporting and Export: Clarity and accessibility of evaluation outputs, with options to export results (e.g., JSON, CSV) and create visualizations such as bar charts or heatmaps.
  4. 4.
    Indic Language Support: Support for Indic and multilingual datasets such as ARC-IN and FLORES. Evaluators assessed the availability of prebuilt resources.
  5. 5.
    Extensibility: Ease of customizing the framework to add new prompt templates, models, datasets, or evaluation metrics.
[…]
@@ 6 Conclusion and Future Work
In this work, we introduced Eka-Eval, a unified and extensible framework designed to streamline the evaluation of LLMs across 35+ diverse benchmarks and 11 key capabilities, spanning nine major evaluation categories-including support for Indic languages. Through a combination of system design, implementation, and user-centered evaluation, Eka-Eval demonstrated high usability and practicality, achieving the highest participant ratings in four out of five evaluation criteria. These results highlight its effectiveness in enabling reproducible, scalable, and inclusive evaluation workflows.
We plan to expand the framework to over 100 tasks, with a focus on underrepresented Indic languages. Future work will categorize benchmarks by language diversity, task difficulty, and domains such as law, healthcare, and governance. We plan to add dynamic task calibration for context length, ambiguity, reasoning complexity, and India-specific knowledge tasks. Priorities include benchmarks for bias detection, hallucination analysis, privacy risks, domain-specific assessments, and adversarial human review. Finally, we will support phase-wise evaluation across pretraining, fine-tuning, and deployment to provide a comprehensive view of model behavior. These enhancements will be included in Eka-Eval v2.0, strengthening its role as the evaluation standard for Indic and other LLMs.
7 Ethics Statement
This research uses publicly available data without personally identifiable information. All datasets and models comply with their terms of use. The work is intended for academic research. Potential misuse or unintended amplification of biases should be carefully considered before deployment.
@@ A.1 Indic benchmark configuration
As demonstrated in Figure 2, a sample configuration used to evaluate the ARC-Challenge-Indic benchmark across 11 Indic languages. It illustrates how task parameters, templates, and dataset references are modularly specified in Eka-Eval.
@@
A.2 CLI Demonstration
The interactive CLI of the Eka-Eval framework is shown below, which guides users through model selection and evaluation setup. Simplifying benchmarking workflows, it is accessible to both researchers and developers.
@@ A.2 CLI Demonstration
As per Figure 3, users are prompted to select high-level task groups (e.g., Reading Comprehension) during CLI setup. This enables fine-grained benchmarking organization and streamlined selection.
@@
Eka-Eval supports local HuggingFace models and API-based models like Sarvam, Gemma, OpenAI, Claude, and Gemini. Users interactively select model source and configuration through CLI. (see Figure 4)
@@
After selecting a task group, users choose specific benchmarks such as SQuAD, BoolQ, or QuAC for focused evaluation within that domain. (see Figure 5)
@@
The CLI displays final benchmark scores for each model in tabular format, including per-task and average scores. Results are also exported as CSV. (see Figure 6)
@@
As per Figure 7) framework allows users to generate multiple types of visualizations: bar charts, heatmaps, radar plots that are based on completed evaluations.
@@
Figure 8 shows performance breakdown across sub-tasks like BoolQ, SQuAD, and QuAC. It provides intuitive insight into strengths and weaknesses of the model.
@@ A.2.1 Prompt Template System
A critical component of Eka-Eval is its sophisticated prompt management system, which handles diverse evaluation paradigms and languages. The framework implements a flexible template system demonstrated through PIQA benchmark prompt 9:
@@
The Prompt template system as shown in Figure 9 supports zero-shot, few-shot, and chain-of-thought prompting strategies, ensuring consistency across evaluation modes and languages. Users can customize prompt strategies and easily configure them in the benchmark configuration file, as shown in Figure 2.
@@ Appendix B Limitations
While Eka-Eval supports a wide range of benchmarks and model backends, it currently lacks a graphical user interface, relying instead on CLI-based workflows. The framework does not yet support vLLMs and has limited support for detailed error analysis or explainability. Additionally, reproducibility may be affected by changes in external datasets or model versions if not explicitly versioned or cached.
Appendix C Acknowledgements
The authors express their gratitude to Ritika Iyer, Nikhil Mishra, and Rajat Kumar Thakur for their contributions in evaluating the framework, reviewing the manuscript, and reporting the results. We also appreciate the valuable suggestions and feedback provided by Aamod Thakur, Prathamesh Shanbhag and Shailesh Panda.